OpenStudy (anonymous):

solution of |x-1|+|x-2|+|x-3|>=6 is??

OpenStudy (anonymous):

@mathmale please help me

OpenStudy (amistre64):

 adding absolute values means we the proper addition of "lines" within special intervals

OpenStudy (amistre64):

im thinking my a and c are redundant at the moment and just the vertex portions are critical for finding changes in direction

OpenStudy (anonymous):

i didnt get u :(

OpenStudy (amistre64):

|x-1|+|x-2|+|x-3| 1 2 3 ... our critical points 1 2 3 <-------------------------------------------> -(x-1) +(x-1) +(x-1) +(x-1) -(x-2) -(x-2) +(x-2) +(x-2) -(x-3) -(x-3) -(x-3) +(x-3) we have lines that add up in each interval like this

OpenStudy (amistre64):

each interval can then be assessed for an overall view of what is happening

OpenStudy (amistre64):

1 2 3 <-------------------------------------------> -(x-1) +(x-1) +(x-1) +(x-1) -(x-2) -(x-2) +(x-2) +(x-2) -(x-3) -(x-3) -(x-3) +(x-3) ----- ----- ----- ------- -3x+6 x+4 x+0 3x-6

OpenStudy (anonymous):

cant v combine the eqn like...|x-1 + x-2 + x-3| ?? sorry m totally stucked wid this mod

OpenStudy (anonymous):

is this correct?? 3x-6 >=6

OpenStudy (amistre64):

we cant combine them since they change at different parts of the domain

OpenStudy (amistre64):

but yes, we can then subtract 6 from each intervals writeup and see if there is a value int he interval that gets us a zero

OpenStudy (anonymous):

and 3x>=12 x>=4 ?

OpenStudy (amistre64):

1 2 3 <-------------------------------------------> -3x+6 x+4 x+0 3x-6 -3x>=0 x-2>=0 x-6>=0 3x-12>=0

OpenStudy (amistre64):

might have to dbl chk my additions for the intervals but thats the basic run of it

OpenStudy (amistre64):

-3x>=0 x-2>=0 x-6>=0 3x-12>=0 x <= 0 x >=2 x>=6 x >=4 good iffy not in good

OpenStudy (amistre64):

-x+4-6 for the second one :) forgot a - -x-2 >= 0 -x >= 2 x <= -2 is not in the intervla